### v1

# Modeling Extensive Defects in Metals through Active Machine Learning and Automated Configuration Reconstruction

Abstract
Extended defects such as dislocation networks and complex grain boundaries are ubiquitous in metals, and accurately modeling these extensive defects is crucial for understanding their deformation mechanisms. Existing machine learning interatomic potentials (MLIPs) often fall short in adequately describing these defects, as their significant characteristic sizes exceed the computational limits of first-principles calculations. To address these challenges, this study constructs a comprehensive defect genome, derived through active machine learning from classical empirical potentials, and augmented by conventional on-the-fly active learning. Additionally, we have developed a approach for transforming defect atomic clusters into periodic configurations via precise atom insertion, using Grand Canonical Monte Carlo simulations. These strategies enable the development of highly accurate and transferable MLIPs for modeling extensive defects in metals. Using body-centered cubic tungsten as a model system, we develop an MLIP that reveals previously undocumented plastic mechanisms in simulations of nanoindentation and dislocation-grain boundary interactions. This framework not only improves the modeling accuracy of extensive defects in crystalline materials but also establishes a robust foundation for further advancement of MLIP development through the strategic use of defect genomes.
Abstract
INTRODUCTION
In the intersecting realms of computational chemistry, materials science, and mechanics, machine learning has made substantial strides, particularly through the creation of machine learning interatomic potentials (MLIPs). These innovative tools have transformed atomic-scale simulations by facilitating the accurate modeling of complex materials behavior with quantum-level accuracy  (?, ?, ?, ?, ?, ?). A diverse array of MLIPs leveraging unique descriptors has emerged. For instance, Neural Network Potentials (NNPs) employ atom-centered symmetry functions (ACSF)  (?), while Gaussian Approximation Potentials (GAP) utilize the smooth overlap of atomic positions (SOAP)  (?), and Spectral Neighbor Analysis Potentials (SNAP) depend on bispectrum components  (?). Additionally, both Moment Tensor Potential (MTP)  (?) and Atomic Cluster Expansion (ACE)  (?) have found extensive application across various materials, from metals and alloys to 2D materials and complex systems like carbon and silicon  (?, ?, ?). A Pareto analysis reveals that while GAP achieves the highest accuracy, it lacks computational efficiency  (?). In contrast, MTP and ACE present a more balanced profile of accuracy and computational demand, with ACE being notably faster and less resource-intensive than MTP  (?, ?, ?).
An essential element in developing MLIPs is constructing a good training database that adequately covers the configuration space relevant to the intended simulations. Typically, this database is curated by domain knowledge, such as ground-state structures, structures deformed under various elastic strains, ab initio molecular dynamics (AIMD) configurations at different temperatures, and defects such as vacancy, interstitial, crack tip and dislocation core  (?). Advancements in machine learning have led to the adoption of active learning (AL)  (?, ?, ?), a more sophisticated approach that selects new configurations efficiently by assessing uncertainty without the need for density functional theory (DFT) calculations. The effectiveness of AL strategies hinges on the ability to calculate similarities among configurations by comparing their (local atomic environments) LAEs. MTP with AL, utilizing the D-optimality criterion and the MaxVol algorithm, exemplifies one of the most widely employed MLIPs, which has been used in numerous studies  (?, ?). Other AL strategies, such as uncertainty-driven dynamics  (?) and hyperactive learning  (?), also significantly contribute to efficiently sampling the configuration space.
Two primary challenges persist in the construction of databases for the development of MLIPs. The first challenge is the comprehensive coverage of defect. For metals, it is relatively straightforward to train MLIPs that can reproduce lattice constants, elastic constants, general stacking fault energy, simple grain boundaries, and dislocation migration. However, it remains unclear whether existing MLIPs can effectively model extended defects and their interactions, even if some of them are deemed for general-purpose  (?, ?, ?, ?, ?). This uncertainty primarily stems from the lack of quantification for unknown defects. Although on-the-fly AL (OTF-AL) has been proposed to accelerate the development of MLIPs, it is insufficient to sample comprehensive defects in every loading condition. Qi et al. suggested that an ideal strategy would be to efficiently generate and sample configuration space before conducting any DFT calculations  (?). They used M3GNet  (?) as the engine for effective and efficient database generation for atomic hydrogen diffusion in titanium hydrides system. This strategy however is unsuitable for modeling extended defects in metals due to the low accuracy and computational efficiency of universal MLIPs in large-scale simulations.
The second challenge involves the accurate modeling of extended defects. For instance, while current MLIPs can handle local defects such as vacancies, screw dislocations, and simple grain boundaries, they often exclude extended defects such as dislocation nucleation and multiplication, general grain boundaries (GBs), and interactions among them  (?). This exclusion is due to the inability of these extended defects to fit within the small periodic configurations typically used in DFT calculations. One workaround has been to extract a non-periodic atomic cluster with a spherical shape from large-scale simulations. To restore periodic boundary conditions required for plane-wave DFT calculations, a sufficiently large vacuum layer is added to prevent image interaction, as implemented in the MLIP-3 package  (?). This strategy has been applied in nanoindentation simulations  (?) and elastic constant calculations of polycrystals  (?) for diamond. Nevertheless, this approach can introduce irrelevant free surfaces or other boundaries  (?). To overcome these limitations, new strategies have been developed to maintain periodic boundary conditions without additional vacuum layers. For instance, Zhang et al. created a periodic crack-tip DFT cell through duplication and rotation operations  (?). Hodapp et al. utilized a screw dislocation in BCC W as an example to construct a periodic configuration  (?), while Mismetti et al. designed periodic configurations containing leading and trailing parts of edge dislocation in face-centered cubic aluminum (FCC Al)  (?). However, these methods fall short in handling aforementioned extended defects.
In this study, we introduce a generalized framework for developing MLIPs specifically designed to model extensive defects in metals. This framework integrates two key components: (1) Empirical Interatomic Potentials-assisted Active Learning (EIP-AL), and (2) Periodic Configuration Construction via Grand Canonical Monte Carlo simulations (PCC-GCMC). These components are crucial for overcoming the challenges previously discussed. The EIP-AL method is designed to enhance defect sampling capabilities during large-scale simulations across a variety of loading conditions. When combined with conventional OTF-AL, this broad sampling is crucial for capturing the full spectrum of defects and their complex interactions, ultimately facilitating the construction of a comprehensive defect genome. On the other hand, the PCC-GCMC technique converts non-periodic atomic clusters of selected defects, into periodic configurations without vacuum. This conversion is essential for ensuring that these configurations are compatible with standard plane-wave DFT calculations. Focusing on body-centered cubic tungsten (BCC W), our analysis underscores the critical role of the EIP-AL and PCC-GCMC techniques in enhancing the predictive accuracy and reliability of MLIPs for large-scale simulations. Ultimately, we have developed a set of MLIPs optimized for BCC W, capable of simulating a wide range of plastic deformations and crack propagation. We demonstrate the capabilities of our MLIPs through nanoindentation simulations on single crystals and compressive tests on spherical bicrystals featuring both low-angle and high-angle twist GBs. These tests reveal new deformation mechanisms not previously documented, underscoring the robustness and adaptability of our MLIPs in handling complex simulation scenarios.
RESULTS
MLIP development framework for modeling extensive defects
In this study, we introduce an generalized framework designed to develop versatile machine learning interatomic potentials (MLIPs) for accurately modeling extensive defects in metals, as depicted in Fig. 1. This framework is implemented through a sequential three-step process. Initially, a comprehensive pool of defect structures is generated using large-scale simulations based on well-established EIPs, notably the Embedded-Atom Method (EAM) and the Modified Embedded-Atom Method (MEAM) for metals. These potentials have been instrumental for over three decades, enabling the rapid simulation of complex defects across millions of atoms, albeit with lesser accuracy compared to DFT. Such capabilities are crucial for simulating a variety of deformation scenarios, including but not limited to general GBs in polycrystals, as well as compression, tension, and shear tests in both polycrystalline and single-crystal systems, and nanoindentation (Fig. 1a). This methodological versatility fosters the creation of a diverse set of local atomic environments for defects (LAEDs), addressing phenomena such as grain boundary relaxation and deformation, dislocation nucleation from material inside and exposed surfaces, dislocation multiplication, and complex dislocation-grain boundary interactions. Subsequently, the D-optimality criterion-base AL implemented in MLIP-3  (?) is used to select representative LAEDs based on the local atomic neighborhoods. This strategy meticulously evaluates the extensive pool of generated configurations against an established basic dataset, enabling the precise selection of the most representative LAEDs. This process yields distinct atomic clusters with typical size of 100-200 atoms. For these clusters, vacuum layers of 8 Å thickness are added in all three dimensions of the simulation box to minimize image interaction. We develop an MLIP, named C-MLIP-0, combining the basic dataset with these cluster configurations. This MLIP is also the result of employing the MLIP-3 approach  (?).
In the second step, our PCC-GCMC strategy is employed to convert all clusters from the first step into periodic configurations as shown in Fig. 1b. For each atomic cluster, we begin by initializing the simulation box size slightly larger than the atomic coordinates in all directions. Subsequently, new atoms (depicted in blue in Fig. 1b) are inserted into the box via GCMC simulations, continuing until no additional atoms can be accommodated. During this insertion phase, the core atoms (shown in red) remain fixed to preserve the original LAE of the center atom. After the insertion is complete, both the coordinates of the newly inserted blue atoms and the box size are relaxed simultaneously. This step is crucial to prevent highly distorted LAEs among the newly inserted atoms. This procedure is iteratively repeated, gradually increasing the box size to facilitate further atom insertions. Through this iterative process, the configuration that exhibits the lowest potential energy throughout all iterations is selected as the final configuration for further single-point DFT calculations. The resulting new DFT dataset, combined with the basic dataset, is utilized to train a foundational model, termed P-MLIP-0, designed to be suitable for simulating common defects. To demonstrate the functionality of our approach, we use the example of a screw dislocation core in BCC W to detail the PCC-GCMC, as described in Supplementary Note 1 and illustrated in Fig. S1. Remarkably, without any manual adjustments, our method successfully generates two distinct periodic configurations: one featuring a screw dislocation dipole and another with a non-compact core in addition to the original compact core. These two periodic configurations are readily prepared for DFT calculations and are instrumental for the development of MLIPs specifically tailored for modeling screw dislocation phenomena.
The final stage, Step 3, is the MLIP refinement phase, as illustrated in Fig. 1c. For specific applications not addressed in the initial EIP-AL of Step 1, such as crack propagation and radiation damage, further EIP-AL and PCC-GCMC operations are required, utilizing appropriate EIPs. A limitation of EIP-AL is its potential bias in generating necessary LAEDs due to the inherent bias of EIPs. Consequently, new LAEDs may be generated during MLIP-based large-scale simulations. To ensure the transferability of MLIPs across diverse applications, we implement the traditional OTF-AL process here for specific cases. All generated atomic clusters are converted into periodic configurations via PCC-GCMC and reincorporated into the training dataset. This process is repeated until no new LAEDs are produced by OTF-AL. The final version of our MLIP, named P-MLIP-1, represents the culmination of our development efforts and is designed to robustly model a wide range of defect phenomena.
Figure 1: The framework of versatile MLIP development for modeling extensive defects. a Empirical interatomic potential-informed active learning (EIP-AL). b Construction of periodic configurations from non-periodic atomic clusters through precise atom insertion using grand canonical Monte Carlo simulations (PCC-GCMC). c MLIP refinement by combining EIP-AL, on-the-fly (OTF-AL), and PCC-GCMC.
Structural analysis of initial defect pool based on empirical interatomic potential-based active learning
This section explores the differences between the cluster dataset shown in Fig. 1a and the periodic dataset from Fig. 1b, alongside the basic dataset informed by domain knowledge. We will examine their localized atomic environments (LAEs) and assess MLIP training performance to identify key distinctions. This examination is conducted using the SOAP descriptor  (?) with parameters set as $r_{\text{cut}}$ = 5 Å, $n_{\text{max}}$ = 12, and $l_{\text{max}}$ = 10  (?). The results are visualized in Fig. 2a, where the first two principal components (PCs) are plotted against each other. In the basic dataset, the majority of atomic environments exhibit PC1 values less than 0, with some data points extending into a long arm, representing surface atoms. The cluster dataset, conversely, divides into two groups: a smaller group that aligns closely with the main portion of the basic dataset and a larger group that positions itself above the extended arm of the basic dataset. This larger group primarily comprises surface atoms, a logical outcome given their proximity to the free surface. The smaller group represents target atoms situated at the center of each cluster configuration. The size disparity between these groups indicates that only a small fraction of LAEs directly contribute to target defects, while the majority, being surface atoms, do not inform the force and energy calculations of non-surface atoms. Unlike the cluster dataset, the periodic dataset forms a single cohesive group containing all target atoms from the cluster dataset. There is an overlapping region between the periodic and basic datasets, marked by a dashed line, suggesting shared common LAEs. This overlap signifies that the periodic configurations created via our PCC-GCMC method exhibit characteristics similar to those identified in Ref  (?), with data points in this region likely corresponding to the perfect BCC lattice. Additionally, other data points in the periodic dataset represent inserted atoms from the GCMC process. Although these LAEDs are not explicitly defined in the initial EIP-AL (such as non-compact screw core discovered in Fig. S1), they may prove instrumental in exploring other types of unknown plastic deformations, which we will discuss further in the following section.
Next, we proceed to train two sets of MLIPs, i.e. C-MLIP-0 and P-MLIP-0, utilizing the cluster dataset (C) and the periodic dataset (P), respectively, each integrated with the basic dataset (B). MTP is used for training. The training performance for these models is assessed by comparing the root mean square error (RMSE) of configuration energies, as presented in Fig. 2(b). It is consistently observed that the B+P combination outperforms B+C across various MTP levels. This superior performance of B+P can be attributed to the more straightforward nature of the dataset, which is devoid of the complex, irrelevant LAEs like surface atoms found in the cluster dataset (illustrated in Fig. 2a). This finding underscores the first significant advantage of employing the periodic dataset in MLIP training: the elimination of irrelevant LAEs enhances the training process, leading to more accurate and reliable MLIPs.
Figure 2: Structural analysis of defect pool from EIP-AL. a Principal component analysis of different datasets. b MTP training performance of two datasets with respect to different levels.
To further elucidate the superiority of PCC-GCMC and the resultant P-MLIP-0, we undertake additional validation tasks by comparing the energy and forces predicted by our MLIP-0 models against new DFT calculations for various GBs relaxed at temperatures of 300 K, 900 K, and 2000 K, as well as under severe compression conditions. The stable structures of these GBs are referenced from an established study in Ref  (?). The GB structures and comparative results for energy and force between MLIPs and DFT are displayed in Fig. S2. Observations reveal that both C-MLIP-0 (trained by B+C) and P-MLIP-0 (trained by B+P) consistently outperform B-MLIP-0 (trained solely by B) across all validation scenarios. An exception is noted in the energy comparison at 2000 K, where B-MLIP-0 records a lower error than C-MLIP-0 and P-MLIP-0 (6.3 vs 12.2 and 7.2 meV/atom). However, the force RMSE for B-MLIP-0 is significantly higher compared to C-MLIP-0 and P-MLIP-0 (199.3 vs 172.2 and 154.4 meV/Å). These findings suggest that incorporating either cluster (C) or periodic (P) datasets enhances the accuracy. Notably, P-MLIP-0 consistently surpasses C-MLIP-0 in both energy and force accuracy, particularly under compression conditions (6.7 vs 10.1 meV/atom for energy; 150.8 vs 173.7 meV/Å for force). This superior performance can be attributed to the inclusion of additional atoms by PCC-GCMC in Fig. 2a, which introduces relevant LAEDs absent in B+C, and thus better aligns with the new validation tasks in Fig. S2. Combining the training performance illustrated in Fig. 2b, we conclude that B+P offers superior training outcomes and higher transferability than B+C, thereby validating its enhanced efficacy in modeling complex defects.
Uncertainty quantification in general plastic deformation based on the initial defect pool
Our MLIP-0 models, C-MLIP-0 and P-MLIP-0, are designed to accommodate a wide range of defect simulations including dislocations, GBs, dislocation nucleation and multiplication, and dislocation-GB interactions. To verify the transferability of our MLIP-0 models, we utilize per-atom uncertainty quantification available in MLIP-3  (?), assessing the extrapolation grade, $\gamma$, to evaluate various defects in BCC W. Values of $\gamma$ ranging from 0 to 1 signify interpolation, whereas values exceeding 1 imply extrapolation. Figs 3a-c display three common types of dislocation loops encountered in the plastic deformation of BCC metals: a $<100>$ interstitial loop, a $<111>$ glide loop, and a $<111>$ climb loop. Given that the performance outcomes of the basic dataset combined with cluster one (B+C) and the basic dataset with periodic one (B+P) are comparable, we present only the results for B+C for clarity. Notably, using the basic dataset (B), all atoms within the $<100>$ interstitial loop and those near the edge segments of the $<111>$ glide loop, along with some atoms in the $<111>$ climb loop, exhibit $\gamma$ values significantly greater than 1, indicating a higher level of uncertainty (left panels of Fig. 3a-c). In contrast, all atoms within these dislocation loops, when analyzed using the B+C dataset, show $\gamma$ values less than 1. This demonstrates our MLIP-0 models’ enhanced effectiveness in capturing complex dislocations, with energies and forces of all LAEs being obtainable through interpolation.
Additionally, we assess the transferability of our models regarding general GBs in polycrystals and their temperature-induced evolution. We calculate the extrapolation grade $\gamma$ for all the atoms in a polycrystal relaxed by a MEAM potential  (?) at 300 K and 2000 K. It should be noted that our EIP-AL in the initial EIP-AL only use EAM-Zhou potential  (?), such that the GBs obtained by MEAM may have different atomic structures. Fig. 3d illustrates the polycrystal, with each atom color-coded according to $\gamma$ values obtained from B+C. Notably, all atoms in this representation exhibit $\gamma$ values less than 1, indicating low uncertainty and high confidence in the predictions. Figs 3e and f display histogram plots of the $\gamma$ values for this polycrystal calculated using different datasets: the basic dataset (B), a dataset from a Gaussian Approximation Potential (GAP)  (?), B+C, and B+P at 300 K and 2000 K. The analysis reveals that both datasets B and GAP predict $\gamma$ values greater than 1 for atoms at both 300 K and 2000 K, indicating higher uncertainty. This is particularly pronounced for B. In contrast, $\gamma$ values calculated by datasets B+C and B+P are consistently below 1 at 300 K, suggesting a higher accuracy in these configurations. However, at 2000 K, while most atoms still exhibit $\gamma$ values below 2, there are a few exceptions with values slightly higher than 2. This indicates that although the models perform well at lower temperatures, at higher temperatures the uncertainty escalates but stays manageable. These results underscore the high transferability and effectiveness of generated defect pool in simulating the behavior of general GB and the thermodynamics of polycrystals.
We further evaluate the uncertainty quantification of our defect pool during a continuous nanoindentation simulation, illustrated in Figs 3(g-i). Figs 3(g, h) display the distribution of defect atoms color-coded by $\gamma$ values at the maximum depth (4 nm), as calculated using B and B+C, respectively. For the case of B, several regions exhibit $\gamma$ values above 1, particularly in regions involving dislocation interactions and surface regions, as indicated by red arrows in Fig. 3(g). In contrast, all defect atoms analyzed with B+C demonstrate $\gamma$ values lower than 1 (Fig. 3(h)), indicating significantly lower uncertainty. Additionally, we calculate the maximum $\gamma$ values ($\gamma_{\text{max}}$) among all defect atoms for each snapshot throughout the nanoindentation process. Comparisons of $\gamma_{\text{max}}$ for B, GAP, B+C, and B+P reveal that B and GAP consistently show $\gamma_{\text{max}}$ values above 2, reflecting high uncertainty throughout the nanoindentation process. Conversely, both B+C and B+P maintain $\gamma$ values around 1, with B+C exhibiting slightly lower $\gamma$ values than B+P. This difference is attributed to B+C having a higher proportion of surface atoms, which are prevalent in open-surface nanoindentation simulations. Nonetheless, these results indicate that our defect pool encompasses all the plastic LAEs during nanoindentation, including dislocation nucleation from open surfaces and complex multiplication beneath the indenter.
Figure 3: Uncertainty quantification of various defects in BCC W based on different dataset. a-c Three types of dislocation loop, $<100>$ interstitial loop, $<111>$ glide loop, and $<111>$ climb loop. Left panels are results from B, and right panels are results from B+C. Blue lines are edge dislocation segments, and red lines are screw segments. d Polycrystal model. e, f The probability density function (PDF) of extrapolation grade ($\gamma$) at 300 K and 2000 K. g, h Defect structures at the maximum indentation depth from B and B+C. i Evolution of the maximum extrapolation grade ($\gamma_{\text{max}}$) during the whole indentation process. All atoms in a, b, c, d, g, h are color-coded by $\gamma$. Red arrows in b, c, g indicate the atoms with high uncertainty. A value of $\gamma$ between 0 and 1 indicates interpolation, while a $\gamma$ greater than 1 suggests extrapolation.
To further highlight the greater diversity of LAED in the B+P dataset compared to B+C, we calculate the extrapolation grade $\gamma$ for configurations within the GAP dataset, evaluated by both B+C and B+P. The comparative results are illustrated in Fig. S3a, where we plot the $\gamma$ values for B+C and B+P against each other. Our analysis reveals that the B+P dataset includes 59 configurations with $\gamma$ values exceeding 5, notably, none exceeding 20. In contrast, the B+C dataset shows 74 configurations with $\gamma$ values over 5, including 8 configurations where $\gamma$ surpasses 20. This stark contrast underscores that B+P offers a richer diversity of LAEs than B+C. Interestingly, Fig. S3a also indicates that certain LAEs present in the GAP dataset are missing in B+P. These absent LAEs are primarily configurations of dumbbell interstitials, which are particularly relevant to simulations of radiation damage. Our current EIP-AL procedure in Fig. 1a does not encompass these scenarios, resulting in their absence. Nonetheless, incorporating these specific LAEs into our model through EIP-AL and PCC-GCMC would be straightforward. Furthermore, we calculate the extrapolation grade $\gamma$ for configurations within the dataset B+P based on the dataset GAP, as depicted in Fig. S3b. It is observed that 14.8% of configurations within B+P are absent in GAP. This significant absence explains the high uncertainty observed for the dataset GAP in Fig. 3, indicating that the existing GAP dataset and the resultant MLIP are not adequately equipped for general defect modeling of BCC W. This finding underscores the necessity of our study and the development of a more comprehensive dataset such as B+P, which better addresses the complexities of large-scale simulations.
MLIP refinement via on-the-fly active learning
It should be noted that all LAEDs discussed above are generated using the EAM-Zhou potential  (?) and the MEAM potential  (?). The applicability of the MLIP-0 models developed based on these LAEDs for modeling extended defects in large-scale simulations, particularly within MLIP simulations, remains uncertain. To address this issue, we employ the traditional on-the-fly AL (OTF-AL) implemented in MLIP-3  (?). This process allows us to monitor the extrapolation grade $\gamma$ during MLIP-based simulations and to identify unknown LAEDs. These activities correspond to Step 3 in Fig. 1c.
Initially, we focus on the high-temperature relaxation of a polycrystal. Although Fig. 3e illustrates that our dataset constructed via the initial EIP-AL encompasses all LAEDs in a polycrystal relaxed by the MEAM potential at 300 K, Fig. 3f reveal that these LAEDs do not adequately represent all LAEDs at 2000 K. Consequently, EIP-AL is deemed ineffective for this high temperature. Here, we engage OTF-AL to accommodate polycrystal relaxation at various temperatures: 300 K, 900 K, 1200 K, and 2000 K, starting with the dataset B+P. In the initial AL loop, no new configurations are generated at 300 K and 900 K. However, the simulation halts at 1200 K due to the loss of atoms, prompting the selection of 21 new configurations with $\gamma$ values slightly greater than 2. These clusters are then transformed into periodic configurations via PCC-GCMC and reintegrated into our training database for the development of an updated MLIP. In the subsequent loop, the relaxation simulations proceed smoothly across all the temperatures. During the 2000 K relaxation, an additional 25 new configurations are identified. These activities ensure that we now possess a comprehensive defect pool necessary for modeling general GBs in a random polycrystal, all of which are embedded in periodic configurations via PCC-GCMC. All these configurations are collectively combined as the dataset OTF-AL-GB. To further asses the uncertainty of our MLIP model trained by B+P+OTF-AL-GB, we use ensemble learning to quantify the uncertainty of a polycrysal relaxed at 300 K and 2000 K (Fig. S4). It is seen that the uncertainty is significantly reduced using the new MLIP models. The details of the ensemble learning are provided in Supplementary Note 2.
We next explore a new application: crack propagation in BCC W across four different crack systems. It is important to note that the initial EIP-AL in Fig. 1a did not include fracture simulations. Consequently, we commence with conducting crack propagation simulations using both EAM-Zhou  (?) and MEAM potentials  (?), known for their wide use in fracture analysis. Critical snapshots of crack propagation are presented in Figs 4a-d using EAM and Figs 4e-h using MEAM. MEAM, specifically tailored for better fracture simulation  (?), demonstrates distinct fracture behaviors compared to EAM. For instance, EAM simulations reveal a BCC-to-FCC phase transition at the crack tip within the (010)[001] crack system (Fig. 4a), while MEAM predominantly shows brittle cleavage (Fig. 4e), although a similar phase transition appears in later stages, possibly influenced by boundary conditions. Additionally, EAM captures dislocation emission in the crack systems $(\bar{1}10)[001]$ and $(\bar{1}10)[\bar{1}\bar{1}1]$ (Figs 4b, d), in contrast to the brittle cleavage observed with MEAM in these systems (Figs 4f, h). Both potentials, however, depict deformation twinning at the crack tip in the crack system $(112)[1\bar{1}0]$ (Figs 4c, g). To refine our MLIPs, EIP-AL is applied to simulation outputs from both EAM and MEAM, resulting in the selection of 14 new defect clusters crucial for fracture simulations. These clusters are then processed using PCC-GCMC to create periodic configurations suitable for further DFT simulations, forming the new dataset, EIP-AL-fracture. Building upon this, a new MTP is trained incorporating the datasets B+P+OTF-AL-GB+EIP-AL-fracture. Subsequent OTF-AL for fracture simulations across the four crack systems leads to the selection of 25 additional configurations. After converting these into periodic configurations via PCC-GCMC and calculating them with DFT, they are integrated into the comprehensive datasets, culminating in B+P+OTF-AL-GB+EIP-AL-fracture+OTF-AL-fracture.
To demonstrate the reduction in uncertainty achieved by integrating the EIP-AL-fracture and OTF-AL-fracture datasets, we apply ensemble learning to calculate the standard deviations of atomic forces for critical snapshots during the OTF-AL of fracture simulations across four crack systems. These calculations are conducted using five independent MTPs developed from datasets B+P+OTF-AL-GB and B+P+OTF-AL-GB+EIP-AL-fracture+OTF-AL-fracture, respectively. The results are displayed in Figs 4i-p. The analysis reveals that, prior to including crack-related LAEDs, there is significant uncertainty at the surfaces of crack tips and in dislocations nucleated from these crack tips. However, after incorporating the EIP-AL-fracture and OTF-AL-fracture datasets, the uncertainty for all atoms is markedly reduced. This improvement is showcased by lower standard deviations in atomic forces, highlighting the effectiveness of the combined EIP-AL and PCC-GCMC approaches in adapting to new applications such as fracture simulations. These findings underscore that the final datasets, enriched with a comprehensive defect genome, offer robust capabilities for simulating general plastic deformations and crack propagation in BCC W, providing a solid foundation for accurate modeling across various loading configurations and boundary conditions.
Finally, we consolidate all datasets in Table S1 and train a series of MTPs at various levels, alongside a complex ACE potential. Given the high accuracy of ACE in training performance and its predictive capability for fundamental properties of BCC W, coupled with its computational efficiency (Table S2), we select ACE as the definitive MLIP model for conducting subsequent large-scale simulations.
Figure 4: EIP-AL and OTF-AL of fracture simulations. a-h Simulation results of four crack systems using EAM and MEAM potentials. i-l Uncertainty quantification of defects near the crack tip based on MLIPs trained with B+P+OTF-AL-GB. m-p Uncertainty quantification of defects near the crack tip based on MLIPs trained with B+P+OTF-AL-GB+EIP-AL-fracture+OTF-AL-fracture. All atoms in a-h are color-coded by common neighbor analysis: blue atoms have a BCC lattice, green atoms have an FCC lattice, and white atoms have an unrecognized lattice. All atoms in i-p are color-coded by the standard deviation of atomic force.
MLIP applications in nanoindentation and bicrystal compression
We initiate the application of our ACE potential by simulating nanoindentation in single crystal W. Although nanoindentation has been extensively studied using EAM potentials  (?, ?, ?), there are relatively few simulations employing MLIPs. This scarcity is likely due to the challenges associated with accounting for dislocation nucleation from open surfaces and handling complex dislocation interactions, which are difficult to simulate directly with DFT. In a recent study, a tabulated GAP potential was used for nanoindentation of single crystal W  (?), but the uncertainty associated with these simulations was not detailed. Meanwhile, an advanced neural network potential was specifically developed for nanoindentation in single crystal Mo  (?). In Figs 3(g-i), we demonstrate that the LAEDs obtained by the initial EIP-AL successfully encompass all the plastic events observed during nanoindentation simulations conducted with EAM. Building on this foundation, we now present the simulation results obtained with our final ACE potential and examine the uncertainty.
The simulation setup is aligned with the parameters described in Ref  (?), depicted in Fig. 5a. The indentation is conducted along the [100] direction, which is free from boundary constraints, while the other two axes maintain periodic boundary conditions. The total system spans dimensions of 38 $\times$ 38 $\times$ 32 $\text{nm}^{3}$, encompassing 2,832,200 atoms. To stabilize the substrate, the bottom layers, with a thickness of 6.4 Å, are fixed. Additionally, several layers, totaling a thickness of 2.56 nm, function as thermostat layers to dissipate heat generated during the nanoindentation process. The remaining atoms are designated as Newtonian layers, where the indentation process occurs. The indentation is performed at a velocity of 20 m/s, using a spherical indenter with a radius of 6 nm. Fig. 5b displays the force-depth curve, reaching a maximum depth of 3.8 nm. The corresponding microstructures at this depth are depicted in Fig. 5c, where all defect atoms are color-coded based on the extrapolation grade ($\gamma$). Notably, all defects exhibit $\gamma$ values lower than 1.2, affirming the low uncertainty in modeling the complex dislocation network generated by the ACE potential. The dislocation pattern reveals prismatic loops advancing along 111 slip directions. This pattern, which includes the process of shear loop formation and cross-slip repeating beneath the indenter tip, aligns with results from previous studies using tabGAP and experimental observations  (?). Additionally, the slip traces and pileup, detailed in Fig. 5d, resonate with findings from the customized neural network potential employed for Mo nanoindentation  (?). These results not only validate the effectiveness of the ACE in capturing intricate deformation mechanisms but also demonstrate its consistency with established models and experimental data.
Our simulations reveal a new mechanism for the onset of plasticity during nanoindentation, as analyzed using BCC Defect Analysis (BDA)  (?). The outcomes of this analysis are illustrated in Fig. 5e, where nano-screw dislocations, deformation twinning, 110 planar faults, surface features, and vacancies are identified. Specifically, the sequence of events leading to the initiation and evolution of plasticity is marked by five critical points A-E, labeled in Fig. 5b and detailed in Fig. 5e. At point A, plasticity begins with the formation of a local hexagonal close-packed (HCP) region. The first yield, at point B, is associated with the formation of three intersecting twin boundaries (TBs). At point C, the vertical TB extends deeper into the material while the left one disappears. At point D, the vertical TB further advances, and the right TB transitions into a regular GB. Finally, at point E, the extension of the vertical TB halts. These results indicate that while deformation twinning is critical at the initial stages of nanoindentation (from B to D), dislocation nucleation and multiplication subsequently take precedence, governing the later stages of deformation (from D to E).
Figure 5: Simulation results of nanoindentation in BCC W. a Simulation setup. b Load-depth curves with five critical points A-E labeled. c The defect pattern at maximum depth, with all atoms color-coded by the extrapolation grade $\gamma$. Arrows indicate the {111} slip directions. d Positions of atoms on the top surface; the slip trace is indicated by a circle. e BCC defect analysis (DBA) at five critical points A-E.
We further apply our ACE potential to investigate dislocation-GB interactions within two distinct spherical bicrystals, each characterized by either a low-angle (GB1) or a high-angle twist GB (GB2). The specific lattice orientations for these bicrystals are illustrated in Fig. 6a. Both bicrystal configurations comprise 258,043 atoms. To facilitate dislocation nucleation and subsequent interactions with the GBs, a rigid plane is displaced downward, compressing the bicrystal. Following each loading increment of 0.25 Å, the system is allowed to relax for 5 ps, achieving a loading velocity of 5 m/s. These simulations are conducted using the NVT ensemble at a temperature of 5 K, ensuring precise control over the thermal conditions and mechanical deformations. Fig. 6b displays the atomic structure of the two distinct GBs. It is observed that the low-angle twist GB is characterized by a dislocation network. The mechanical responses of the two bicrystals, labeled BC1 and BC2, under compression are meticulously documented through various metrics in Figs S5 and Fig. 6c. These figures display the contact area-displacement curves, force-displacement curves, and stress-displacement curves, respectively. Analysis of these curves reveals four distinct behavioral regimes, S1 through S4, each demonstrating unique characteristics between the two bicrystals. In particular, in regimes S1 and S2, BC1 and BC2 exhibit comparable contact areas, as shown in Fig. S5a. However, in regimes S3 and S4, BC2 presents a larger contact area. Concurrently, Fig. S5b illustrates that in regime S1, the forces exerted by BC1 and BC2 are equivalent, whereas in regimes S2 and S4, BC2 exerts greater force than BC1, and conversely, in regime S3, BC1 surpasses BC2 in force. The stress, defined as the ratio of force to contact area, is depicted in Fig. 6c. Here, BC1 and BC2 display comparable stress in regimes S1 and S4. BC2 exhibits greater stress in regime S2, while BC1 shows higher stress than BC2 in regime S3. Furthermore, we have computed the maximum extrapolation grade ($\gamma_{\text{max}}$) across the entire compression process, detailed in Fig. S5c. This assessment confirms the robustness of our ACE model in accurately predicting the forces and energy throughout bicrystal deformation, even under extreme strain conditions.
We conduct atomic shear strain analysis and extract dislocation lines to understand the distinct deformation mechanisms in the two bicrystals, BC1 and BC2. The detailed analyses of atomic slip histories for BC1 and BC2 at four critical points A-D, as labeled in Fig. 6c, are presented in Figs 6d ,e. Fig. S6 displays the extracted dislocation lines within each bicrystal. In the initial stage, regime S1, both bicrystals exhibit overlapping curves in Fig. 6c and Figs S5a, b due to similar behaviors characterized by elastic deformation and dislocation nucleation, which are independent of GB type. In regime S2, BC2 experiences higher forces and stress compared to BC1. This difference is attributed to the ease of dislocation transmission across GB1, altering its dislocation network, as shown in Fig. 6d-A and Fig. S6a-A. Conversely, all dislocations are impeded by GB2, as depicted in Fig. 6e-A and Fig. S6b-A, demonstrating the influence of GB strengthening or the Hall-Petch effect, which contributes to the increased force and stress in BC2. In regime S3, BC1 exhibits higher force and stress than BC2. This is due to the smaller contact area in BC1 compared to BC2 (Fig. S5a). In BC1, deformation is uniformly distributed across the bicrystal (Fig. 6d-B and Fig. S6a-B), whereas in BC2, deformation concentrates in the upper grain (Fig. 6e-B and Fig. S6b-B), leading to a larger contact area and subsequently lower nucleation stress for dislocations in BC2. In regime S4, the stresses of the two bicrystals become comparable. This results from a balance between GB strengthening and nucleation weakening in regime S4 in BC2, as illustrated in Fig. 6d-C and Fig. S6-C. In fact, BC1 behaves similarly to a single crystal throughout the compression process because the low-angle twist GB exhibits a negligible blocking effect on dislocation movement (Fig. 6d-C and Fig. S6a-C). Conversely, in BC2, the high-angle twist GB acts as a robust barrier, significantly impeding dislocation mobility. Even at the end of the compression process (point D), this effect is evident as the upper grain of BC2 is compressed into a flat layer, while the lower grain retains its original spherical shape, as depicted in Fig. 6d-D and Fig. S6b-D. Moreover, significant GB sliding is observed in BC2, particularly highlighted in Figs 6c-C, D. These findings underscore the capability of our ACE model to accurately simulate and analyze complex plastic deformations involving intricate interplays between dislocation motion and GB deformation.
Figure 6: Simulation results of bicrystal compression. a Simulation setup and lattice orientations of upper and lower grains. b GB structures. c Stress-displacement curves of two bicrystals. Four regimes S1-S4 and four critical points A-D are identified. Atomic slip history of the bicrystal d containing a low-angle GB and e containing a high-angle GB. All atoms in d and e are color-coded by the atomic shear strain.
DISCUSSION
Over the past decade, MLIPs have emerged as a cornerstone in computational science, significantly impacting fields such as chemistry, materials science, and mechanics  (?, ?, ?, ?, ?, ?). Various frameworks employing a range of atomic descriptors, including ASCF  (?), SNAP  (?), moment tensor  (?), and ACE  (?), have been developed. These frameworks have implemented diverse training strategies, from linear regression to neural networks, primarily focusing on enhancing the accuracy and computational efficiency of MLIPs. Such advancements are crucial, as they significantly improve the ability of MLIPs to model complex atomic interactions both accurately and efficiently, thereby paving the way for groundbreaking discoveries in multiple scientific domains. However, the construction of a robust training dataset is equally essential. The quality and comprehensiveness of the data used for training directly impact the effectiveness and reliability of MLIPs. Despite these developments, the scrutiny applied to most MLIPs remains insufficient. These potentials are often deployed in large-scale applications without undergoing rigorous validation across diverse conditions or against a wide spectrum of atomic configurations. Consequently, due to inadequately comprehensive training datasets, many existing MLIPs fail to accurately model general defects, thereby limiting their applicability in complex simulations.
While the MLIP-3 package facilitates OTF-AL using MTPs to tackle similar challenges  (?), our integration of EIP-AL and PCC-GCMC offers substantial improvements. EIP-AL significantly reduces the demanding requirements of OTF-AL, expediting the development of MLIPs. Typically, OTF-AL involves iterative MD simulations based on MLIPs and frequent retraining  (?), a process that is time-consuming, especially for rare events like diffusion and phase transitions. Moreover, retraining MLIPs—whether through linear regression or neural networks—is costly and computationally intensive. The data in Table S1 highlights this efficiency. Of the 348 configurations generated, only 72 are from OTF-AL. On the other hand, the new LAEDs acquired from OTF-AL are usually specific to the simulated loading conditions and may not generalize well across different scenarios. It is impractical to perform OTF-AL for every conceivable loading condition. For example, certain LAEDs that occur in low strain-rate MD simulations might not emerge under high strain rates, such as the rate-dependent ductile-to-brittle failure transition observed in metal nanowires  (?, ?). Consequently, OTF-AL from MD simulations at high strain rates cannot be used to develop MLIPs suitable for lower strain rates, as critical LAEDs at lower rates remain unexplored, as evidenced in. This strain rate limitation has been identified as a significant challenge in advancing MLIP development  (?). EIP-AL effectively addresses this challenge by leveraging cost-effective EIPs to conduct extensive MD simulations. Even within the same material, various EIPs exhibit distinct properties, which can produce a diverse range of LAEDs under identical loading conditions, as demonstrated by the crack propagation scenarios shown in Fig. 4. For tungsten alone, there are 32 different EIPs available from OpenKIM. The varied characteristics of these EIPs allow us to discover a more extensive array of LAEDs than those typically identified through OTF-AL.
Our PCC-GCMC method significantly refines our approach to developing MLIPs for modeling extensive defects, serving as a critical supplement to both EIP-AL and OTF-AL. This method departs from the conventional approach used in the MLIP-3 package  (?), which involves adding a vacuum layer to minimize image interactions—a process that proves inefficient due to the limited use of LAEs, with most being near the surface (Fig. 2a). Additionally, this traditional approach requires further DFT calculations for each LAED using the Gamma point, a necessity due to the increased box size following the introduction of the vacuum layer. This complicates the training process because of the resulting inconsistent Kpoint density (Fig. 2b). In contrast, PCC-GCMC streamlines the creation of periodic configurations that are immediately integrated into the training dataset following DFT calculations, enhancing the dataset’s relevance, as discussed in Fig. 2 and Fig. S2. Specifically, This method avoids the use of a vacuum layer, circumventing issues with surface atoms. By selecting only the lowest-energy configurations from PCC-GCMC, we ensure that all LAEs in these configurations are physically pertinent to the central LAED. The introduction of atoms in this process could yield perfect lattice or related defect structures (Fig. S1), or more likely, identify previously unrecognized defects that improve the transferability of the developed MLIP (Fig. S2).
Finally, we propose that our framework can be generalized across a diverse array of common crystalline materials to model extensive defects, as illustrated in Fig. 7. This framework is versatile, applicable not only to familiar structures such as FCC, BCC, HCP, and diamond cubic but also to more complex lattices like perovskite. A rich assortment of EIPs is available from resources like the NIST Interatomic Potentials Repository (NIST-IPR)  (?) and OpenKIM  (?). These resources provide access to a variety of potentials, including Embedded Atom Model (EAM), Modified Embedded Atom Method (MEAM), bond order potentials (BOP), Stillinger-Weber (SW) potentials, Tersoff potentials, and reactive force fields (ReaxFF), among others. OpenKIM, for instance, hosts over 644 potentials, with a significant majority being EIPs. Computational materials science has benefited greatly over the past three decades by leveraging these EIPs  (?), and despite the current emphasis on MLIPs, the value of EIPs remains substantial. Although these potentials exhibit higher force and energy errors compared to DFT calculations, their consistent internal logic as physical models provides significant value. Utilizing the abundant EIPs available, our study’s approach can be used to develops advanced MLIPs through three strategic initiatives: adapting EIPs from different materials but sharing the same lattice structure by merely rescaling lattice constants to generate relevant LAEDs; parameterizing EIPs to customize the mechanical properties of materials  (?, ?); and employing the average-atom interatomic potential for random alloys to produce an unlimited series of EIPs from basic elemental forms  (?). Utilizing these resources, our approach enables the creation of a standard defect genome for each crystalline structure. These genomes form the cornerstone for developing MLIPs capable of accurately modeling any type of defect. This not only facilitates further research within this study but also provides a valuable tool for the broader scientific community, enhancing the predictive capabilities of computational material science.
Figure 7: Implications of this study on MLIP development for modeling extensive defects in general crystalline materials.
In summary, this work introduces an advanced framework for developing MLIPs by integrating EIP-AL, PCC-GCMC, and OTF-AL. This framework enables automated, ab initio-accurate, large-scale atomistic modeling of general defects in metals. The developed MLIPs are highly transferable and are based on comprehensive datasets that facilitate simulations of complex behaviors, including general GBs in polycrystals, dislocation nucleation and interactions, dislocation-GB interactions, and crack propagation in BCC W. Rigorous uncertainty quantification analyses, employing D-optimality criterion and ensemble learning techniques, confirm the accuracy of our models in these scenarios. We have successfully applied our trained ACE potential to uncover dominant deformation mechanisms in nanoindentation, identifying deformation twinning as the primary mechanism at the onset of plasticity, and dislocation evolution as the predominant mechanism at later stages. Additionally, we have elucidated the distinct strengthening mechanisms between low-angle and high-angle twist GBs in spherical bicrystal compression tests. This approach proves versatile enough to be generalized for modeling extensive defects in other BCC metals, FCC metals, and high-entropy alloys, demonstrating its broad applicability across various material systems.
MATERIALS AND METHODS
First principles calculations
[…]
Empirical interatomic potentials
In our simulations, we utilize both the Embedded-Atom Model (EAM) and the Modified Embedded-Atom Model (MEAM) to perform large-scale molecular dynamics simulations of body-centered cubic tungsten (BCC W) under various loading conditions. Specifically, we employ the EAM-Zhou potential  (?) for simulating the relaxation, compression, and tension of W polycrystals, as well as single-crystal compression, tension, and nanoindentation processes. Additionally, the MEAM potential referenced in Ref  (?) is applied to model crack propagation in single-crystal W.
Atomistic simulations and analysis
We employ the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) package for all the atomistic simulations  (?). For visualizing atomic configurations and post-processing results such as extracting dislocation structures, we use OVITO  (?). Additionally, Atomsk is used to generate polycrystalline structures  (?). These packages provide a comprehensive suite of tools to support our study of the mechanical properties and plastic deformation mechanisms of BCC W.
Machine learning potential development
We employ the MLIP-2 package  (?) and pacemarker  (?) to develop MTPs and ACE potential for BCC W. MLIP-2 utilizes moment tensor descriptors and apply linear regression to train the machine learning model, enabling it to predict the energy, force, and stress of atomistic systems. Within our study, the weights of energy, force and stress are set as 1, 0.01 and 0, respectively. For ACE, we utilize the highly nonlinear potential energy function with the complex form of $E_{i}=\varphi+\sqrt{\varphi}+\sum_{i}\varphi^{f_{i}}$, where $E_{i}=\varphi+\sqrt{\varphi}+\sum_{i}\varphi^{f_{i}},\quad f_{i}\in\left\{\frac{1}{8},\frac{1}{4},\frac{3}{8},\frac{3}{4},\frac{7}{8},2\right\}$, obtained from Ref  (?). For the expansion of the atomic properties, we used 72 basis functions with 801 parameters. As radial basis we employ Bessel functions. A $\kappa$ value of 0.01, which gives the ratio between force and energy weights value, was used during fitting. For optimisation we use the BFGS algorithm for 2000 steps. The cutoff distance is set as 5 Å for both MTP and ACE. More importantly, we utilize MLIP-3  (?) to perform active learning based on the MaxVol algorithm from extensive simulations conducted with EAM or MEAM potentials. This method was originally developed for the OFT-AL of atomic environments, incorporating a cycle of MTP-based simulations, DFT calculations, and subsequent MTP retraining. In this study, to expedite the process, we exclusively utilize the EIP-AL to identify representative defects from a comprehensive pool of large-scale simulations. This approach and its implications are further elaborated in Section 3.
Data availability
The data that support the findings of this study are available upon reasonable request. The parameter files of the potential and training dataset will be published upon the publication of this paper.
Code availability
All simulations are executed using open-source software LAMMPS. The machine learning force field was trained and validated by the MLIP package (?).
Acknowledgments
This work was sponsored by Nederlandse Organisatie voor WetenschappelijkOnderzoek (The Netherlands Organization for Scientific Research, NWO) domain Science for the use of supercomputer facilities. The authors also acknowledge the use of DelftBlue supercomputer, provided by Delft High Performance Computing Center (https://www.tudelft.nl/dhpc).
Author Contributions
F.S. performed DFT calculations, developed the machine learning force field, performed the atomistic simulations, and wrote the first draft of the manuscript, K.L. and F.S. proposed the PCC-GCMC approach. F.S., L.L. and P.D. conceptualized the project and designed the research. F.S., K.L., Y.J., G.W., L.L. and P.D. analyzed the simulation data. All the authors contributed to the interpretation of the data.
Conflict of Interest
The authors declare no conflict of interest.
Supplementary Materials for
Modeling Extensive Defects in Metals through Active Machine Learning and Automated Configuration Reconstruction
Fei Shuang, Kai Liu, Yucheng Ji, Wei Gao, Luca Laurenti, Poulumi Dey
Corresponding author. P.dey@tudelft.nl; F.Shuang@tudelft.nl.
This PDF file includes:
Supplementary Note 1-3
Table S1 and S2
Figure S1 to S6
Supplementary Note 1. PCC-GCMC for a screw dislocation core
Fig. S1 utilizes the example of a screw dislocation core in BCC W to demonstrate the PCC-GCMC process. As shown in Fig. S1a, the atomic cluster includes a screw dislocation core with three compact core atoms, labeled 1-3. This configuration maintains periodic boundary conditions along the z-direction, with free boundary conditions along the $x$ and $y$ directions. The PCC-GCMC process is repeated with varying box dimensions, $\Delta_{x}$ and $\Delta_{y}$, from 1 to 10 Å. Fig. S1b displays the minimum potential energies per atom post-PCC-GCMC, indicating that the resulting periodic configurations exhibit varied potential energies. Notably, configurations with potential energies below -8.8 eV/atom, depicted in blue, approach the cohesive energy of -8.9 eV/atom and are nearly equivalent to a perfect BCC lattice. Fig. S1c highlights two representative configurations. In the top panel, we observe a periodic configuration featuring a new non-compact core crossing the periodic boundary along the $y$-axis dislocation, with the remainder of the atoms conforming to the perfect BCC lattice structure, excluding the original screw core. The bottom panel reveals a periodic configuration that includes a screw dislocation dipole, akin to typical screw dislocation dipole configurations with a tilted box. Both configurations depicted in Fig. S1c are generated entirely through automated processes. All newly inserted atoms either form a pure BCC lattice or constitute relevant LAEs, making them suitable for direct inclusion in our MLIP training.
Supplementary Note 2. Uncertainty quantification
We adopt two methodologies for uncertainty quantification. The first utilizes atomic descriptors, specifically the extrapolation grade $\gamma$, as implemented in MLIP-3  (?) and similarly applied in ACE  (?), based on the D-optimality criterion and the MaxVol algorithm. A value of $\gamma$ between 0 and 1 indicates interpolation, while a $\gamma$ greater than 1 suggests extrapolation. The second employs ensemble learning, where we train five independent MTPs on the same dataset to compute the standard deviation of atomic force:
$\sigma_{F}=\sqrt{\sum_{i\in\{x,y,z\}}\sigma(F_{\alpha,i}^{(1..N)})^{2}}.$ (S1)
Here, $F_{\alpha,i}^{(1..5)}$ is the $i$-th component of the force acting on atom $\alpha$ as obtained by one of the 5 committee members (MLIPs) for a particular configuration; $\sigma$ denotes the standard deviation. IIt should be noted that while both approaches are valuable for uncertainty quantification, they fulfill different objectives. Each method can identify new local atomic environments (LAEs), but the extrapolation grade, $\gamma$, is specifically designed for this purpose and is particularly useful for on-the-fly active learning. However, merely having a good sampling of LAE does not ensure the effectiveness of an MLIP. It is essential to implement a robust MLIP framework that can effectively train these datasets. Therefore, we integrate both methods in our work to leverage their distinct advantages.
Fig. S4 showcases the uncertainty quantification results of a polycrystal relaxed at 300 K and 2000 K, based on the collective datasets of B+P and on-the-fly AL for GBs (OTF-AL-GB) sessions to train five independent MTPs. We then calculate the standard deviation of atomic forces ($\sigma_{F}$) for all atoms according to Equation S1. Comparative calculations are also performed using five independent MTPs trained using only the basic dataset (B). Figs S4a, c show the GB atoms color-coded by $\sigma_{F}$ at 300 K and 2000 K based on dataset B and the combined datasets B+P+OTF-AL-GB, respectively. Notably, a significant number of GB atoms exhibit $\sigma_{F}$ values higher than 0.5 eV/Å at both temperatures when relying solely on dataset B. In contrast, all atoms show $\sigma_{F}$ values below 0.5 eV/Å at both the temperatures when assessed with the combined datasets. The histogram plots in Figs S4b, d further validate that the MLIPs trained on the collective datasets (B+P+OTF-AL-GB) can accurately predict the atomic forces for all GBs at both 300 K and 2000 K, highlighting the enhanced precision and reliability of our model in high-temperature conditions.
Supplementary Note 3. Final MLIP training and selection
We possess a comprehensive final database that aggregates representative LAEDs in periodic configurations: B+P+OTF-AL-GB+EIP-AL-fracture+OTF-AL-fracture. Detailed breakdowns of each dataset are presented in Table S1. Altogether, the database encompasses 1026 configurations and 94,516 atoms. Of the 348 configurations newly incorporated, 276 are obtained through EIP-AL, while only 72 results from on-the-fly AL. This distribution highlights the efficiency of EIP-AL in significantly reducing the reliance on the more resource-intensive OTF-AL, thereby expediting the development of MLIPs.
The MLIPs trained from this database are equipped to simulate nearly all plastic deformations and fracture of BCC W. However, configurations necessary for radiation simulations due to cascade displacement are currently not included. These configurations could easily be integrated from existing databases or generated through EIP-AL and OTF-AL in future enhancements. We utilize different levels of MTP and a complex ACE potential for training various MLIPs. We particularly favor ACE due to its computational efficiency and the ease of calculating extrapolation grades during simulations in LAMMPS, which greatly facilitates the ongoing assessment. The analysis of basic properties of BCC W, computational efficiency, and training/test performance are detailed in Table S2. It is observed that the accuracy of MTP models increases with their complexity level, but at the cost of reduced computational efficiency. Specifically, the computational expense of MTP20 is seven times that of MTP12. In contrast, ACE demonstrates accuracy comparable to MTP20 while maintaining efficiency similar to MTP12. Based on these findings, ACE is selected as the optimal model for modeling general defects in BCC W due to its superior balance of accuracy and efficiency.
Table S1: Datasets generated in this work at different stages.
$N_{\text{cfg}}$ Average atom number
Basic 678 72
Initial EIP-AL 262 133
OTF-AL-GB 47 135
EIP-AL-fracture 14 118
OTF-AL-fracture 25 115
Total 1026 92
Table S2: Comparison of MLIPs and DFT, and training/test errors of MLIPs. The CPU cost is quantified with respect to EAM for each MLIP.
$a_{0}$ $C_{11}$ $C_{12}$ $C_{44}$ $\gamma_{110}$ CPU cost $E_{\text{rmse}}$ $F_{\text{rmse}}$
(Å) (GPa) (GPa) (GPa) (mJ/m2) (meV/atom) (meV/Å)
DFT 3.184 520.35 199.88 142.42 1772.74
EAM 3.165 519.93 202.17 158.96 1725.98 1
MTP12 3.188 461.87 196.31 116.73 1670.87 28 11.63/11.51 212.01/199.51
MTP14 3.194 480.47 204.37 109.69 1673.98 48 8.66/6.95 196.93/185.71
MTP16 3.189 462.43 185.49 127.46 1630.25 78 7.61/6.31 182.3/171.84
MTP18 3.187 537.04 214.61 140.45 1644.52 125 5.60/4.69 155.71/149.05
MTP20 3.190 530.23 195.56 146.15 1628.74 195 5.02/4.13 141.86/136.46
ACE 3.190 508.05 201.09 139.57 1641.21 30 4.08/4.28 126.26/160.73
Figure S1: Illustration of PCC-GCMC using a screw dislocation core. a Atomic cluster showcasing a screw dislocation core; atoms are color-coded based on Common Neighbor Analysis (CNA) shown in the bottom panel. b Average potential energy per atom for each configuration post atom insertion. c Two representative periodic configurations corresponding to points A and B in b.
Figure S2: Direct validation of MLIPs in nice grain boundaries (GBs) by comparing energy and force from DFT calculations. a Pristine GB structures. b Compressed GB structures. All atoms in a and b are color-coded by common neighbor analysis. c Energy comparison. d Atomic force comparison.
Figure S3: Uncertainty quantification based on the extrapolation grade $\gamma$ for the datasets B+C and B+P. a Comparison of $\gamma_{\text{max}}$ for the GAP dataset obtained based on B+P and B+C. b $\gamma_{\text{max}}$ for B+P based on the GAP dataset. A value of $\gamma$ between 0 and 1 indicates interpolation, while a $\gamma$ greater than 1 suggests extrapolation.
Figure S4: Uncertainty quantification based on ensemble learning for a polycrystal. a Results from the basic dataset. b The probability density function of $\sigma_{F}$ at temperatures of 300 K and 2000 K, derived from the basic dataset. c Results from the combined dataset (B+P+OTF-GB). d The probability density function of $\sigma_{F}$ at 300 K and 2000 K based on the combined dataset.
Figure S5: Additional plots of two crystals compression. a The evolution of contact area with respect to displacement for two bicrystals. b The evolution of force with respect to displacement for two bicrystals. c The evolution of the maximum extrapolation grade ($\gamma_{\text{max}}$) with respect to displacement for two bicrystals.
Figure S6: Extracted dislocation lines during bicrystal compression simulations. A-D correspond to the critical points in Fig. 6c. Left panels are snapshots of bicrystal containing a low-angle twist GB, and right panels are snapshots of the bicrystal containing a high-angle twist GB.